Sol Summary Report – February 15th


Crew 275 Sol Summary Report 15Feb2023

Sol: 3

Summary Title: Waiting for a Martian spring.

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Another very busy day starting with a health status check at 6:45, followed by the daily physical activity session organized by our Health & Safety Officer. We managed to solve a few technical problems during some tests of the Echofinder experiments, whose purpose is to allow non-experts to perform ultrasound measurements using artificial intelligence and augmented reality. Our crew scientist spent a lot of energy reorganizing the schedule due to the snow and the cold temperatures. The crew engineer continued to work on the deployment of environmental sensors that will measure humidity, temperature, and luminosity from different places in the station. But the most important task of the day was certainly the preparation of tomorrow’s EVA, with the labeling of the different parts of the MegaAres instrument and an extensive rehearsal to build it, performed in the lower deck of the Hab.

Look Ahead Plan:

With the snow slowly melting, we decided to keep tomorrow’s EVA dedicated to the deployment of the MegaAres instrument, meant to measure the electric field in the atmosphere. However, we postponed it to the afternoon, so that the temperature is more compatible with the planned activities. If no additional precipitation occurs until Friday, we will go ahead with the deployment of the meteorological mast, which will be necessary for the contextualization of the data from the two atmospheric instruments Mega-Ares and LOAC.

Anomalies in work: Nothing to report.

Weather: Cloudy, low -3°C/27°F, high 1°C/33°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, Operations Report, Journalist Report, HSO Report, GreenHab Report, Astronomy Report.

Support Requested: None.

GreenHab Report – February 15th



GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 70.55 °F

Average humidity: 19%

Hours of supplemental light: 4 hours

Daily water usage for crops: 8.77 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 154.20 gallons

Time(s) of watering for crops: 9:35 & 15:25

Changes to crops: N/A

Narrative: Last night the supplemental light didn’t light up at 10pm, I went to check it out, the timer just had a few minutes offset. I, thus, changed it back to normal; it worked as intended.

Today I maintained and watered crops and plants as my daily operations. There were some new dry leaves on tomatoes, thus, I cleaned them and watered well the tomatoes.

As a new habit, to help their leaves not to be dry, I pulverize water on crops’ leaves to simulate morning dew.

I then ran some water tests on the aquaponics to ensure the safety of the fishes, everything is well.

I also harvested tomatoes and lemon balm at 10:57; it was a wonderful surprise to discover that many tomatoes were ready to be harvested.

Harvest: 177g of tomatoes & 1g of lemon balm

Support/supplies needed: N/A

Operations Report – February 15th


Crew 275 Operations Report 15Feb2023
SOL: 3

Name of person filing report: Quentin ROYER
Non-nominal systems:
• Toilets
• RAM: 1 leak in the roof
• Science Dome: 3 leaks in the walls
• Science Dome: Heater possibly non-nominal during the night
• GreenHab: supplemental light

Notes on non-nominal systems:
• RAM: 1 leak in the roof, probably due to snow. There was no leak in the morning due to freezing temperatures, but this afternoon the bucket which had been placed under the leak was full and had to be replaced.
• Science Dome: As yesterday, 1 leak coming from the junction between the ground and the wall under the window. A small amount of water leaked from the outside and flowed into the Science Dome. The puddle was about 1 cm high, 5 cm wide, and 3 meters long at 3:00 pm, and is located close to an electric socket. As yesterday, 1 leak coming from the junction between the ground and the wall near the door. The puddle was about 0.5 cm high, 3 cm wide, and 70 cm long at 3:00 pm. This second puddle fully dried during the afternoon. A third leak appeared at the junction between the ground and the wall under the window, but in the right side of the window. At 5:00 pm, the puddle was about 0.5 cm high, 4 cm wide and 1.5 meters long, and it is located close to an electric socket.
• Science Dome: as per protocol of an experiment, the Crew Engineer placed an environmental sensor in the Science Dome, including a temperature sensor. A recording of the temperature was performed during the night, and it shows that the temperature decreased during the whole night. Therefore, we assume that the heater did not turn on properly. I can send you the graph of the temperature between 11:55 pm and 6:55 am if you need it.
• GreenHab: supplemental light did not switch on yesterday at 10:00 pm. The Crew Botanist went in the GreenHab to set up the timer at the right time (it was 10 to 15 minutes late).

Spirit rover used: No
• Hours: 208.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: No
• Hours: 220.8
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.4
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: No EVA today – the rovers were not used.

Summary of Hab operations: The crew conducted routine activities, including rehearsal of instruments’ deployment.
• WATER USE: 15.96 gallons
• Water: 424.35 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. Internet for the AI4U experiment (MDRSCampus) was tested today within the boundaries of the protocol of this experiment and is fully operational.

Summary of suits and radios: Spacesuits and radio were not used today, as there was no EVA.

Summary of GreenHab operations: The Crew botanist conducted routine operations and water tests within the protocol of the aquaponics experiment.
• WATER USE: 8.77 gallons
• Heater: On
• Supplemental light: From 10:15 pm to 2:00 am – Timer settings were changed back to nominal
• Harvest: 177 g of tomatoes – 1 g of lemon balm

Summary of ScienceDome operations: The Commander and Crew Scientist tested some experimental protocols.
• Dual split: Heat / From 10 pm to 7 am – We suspect that the dual split did not work this night

Summary of RAM operations: Trainings for the MegaAres experiment were performed. The Crew Engineer repaired a support arm used for the EchoFinder experiment. Environmental sensors were prepared for future deployment.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: There was a small safety issue in the RAM using a welding iron. As the welding iron was still hot, it made contact with a small piece of aluminum which started to melt. The resulting hole is very small, about a centimeter in diameter. He reacted quickly by removing the welding iron when the first vapors arrived, and putting it in the RAM airlock. He contacted the Hab to inform us before turning on the supplemental ventilation of the RAM, as there was no immediate risk for his health.

Questions, concerns, and requests to Mission Support: None

Astronomy Report – February 15th



Astronomy Report

Name: Alexandre VINAS

Crew: 275
Date: 02/15/2023


Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF

Objects Viewed: /

Problems Encountered: snow, wind, and clouds


Solar Features Observed: /

Problems Encountered: snow and clouds…

Journalist Report – February 15th



Sol 3 – Waiting for a Martian spring;

“It was quiet in the deep morning of Mars, as quiet as a cool and black well, with stars shining in the canal waters, […] the moons gone, the torches cold, the stone amphitheaters deserted.”

– Chapter 3 of The Martian Chronicles by Ray Bradbury

Me and my fellow Martians awoke this morning, some of us not as well rested as they hoped, and looked out the Hab window to find that the sky had not cleared, possibly impacting the EVA planned on Sol 4 to install MegaAres and LOAC. After the daily health status check session and a forcibly shortened workout session, we went to work: Corentin, Alexandre and I practiced building the Mega Ares antenna, with a few mishaps and unexpected breakage. Meanwhile, Alice made our first batch of bread (!) while Quentin started mounting the environmental sensors.

The test session for the EchoFinder experiment, designed to aid astronauts in performing ultrasounds using an AI and augmented reality, was a bit chaotic: The support arm used to maintain the tablet was not usable, and one of the two QR cubes used to track the device was wrongly positioned. We will make some adjustments before the first actual session takes place later this week.

After a very active morning, eating Alice’s freshly baked bread for lunch was the perfect fixer upper. A few laughs and a good meal gave us the energy to continue plowing through the day. In the afternoon, the Hab went somewhat quiet as everyone went their separate ways to complete their tasks, and took turns completing cognitive tests and posture analyses. By 6pm, thanks to Alexandre, the MegaAres antenna was ready for tomorrow’s important EVA!

In sum, a quiet Sol, in anticipation and preparation of the next.

Copyright © The Mars Society. All rights reserved. | Main Site